What is the eighth term of the geometric sequence whose first three terms are
3, 6, and 12 Thanks

Answer by Tutorteddy.Com(12)

nth term Tn = ar^(n-1), where a is first term, Tn is nth term and r is the common ratio.
Here, r = 6/3 = 2.
Hence, 8th term = 3 X 2^(8 - 1) = 3 X 2^7 = 3 X 128 = 384